OpenStudy (anonymous):

find the value of x=1 X = 1 + ____1______ . 2+ ____1______ . 2+ ____1______ . 2+ ____1______ . 2+ …….infinity

OpenStudy (anonymous):

It's already the value. :)

OpenStudy (anonymous):

sorry I am just try to give it shape.. not sure how to put underline.. so ...

OpenStudy (anonymous):

I really can't understand the input.

OpenStudy (mathmath333):

is this your question ? \(\large \color{black}{\begin{align} x = 1 + \cfrac{1}{2 + \cfrac{1}{2 + \cfrac{1}{2 + \cfrac{1}{2 + \cdots \infty^{+}}}}} \end{align}}\)

OpenStudy (anonymous):

whew!! mathmath333 you have always been my saviour so far

OpenStudy (anonymous):

@mathmath 333; I don't know how you are able to write like this but yes this is indeed the question Sir,,

ganeshie8 (ganeshie8):

\[\large x = 1+ \dfrac{1}{1+x}\]

OpenStudy (mathmath333):

ganeshie8's method came from this \(\large \color{black}{\begin{align} x = 1 + \cfrac{1}{2 + \cfrac{1}{2 + \cfrac{1}{2 + \cfrac{1}{2 + \cdots \infty^{+}}}}}\hspace{1.5em}\\~\\ \implies x +1= 2 + \cfrac{1}{2 + \cfrac{1}{2 + \cfrac{1}{2 + \cfrac{1}{2 + \cdots \infty^{+}}}}}\hspace{1.5em}\\~\\ \implies x +1= 2 + \cfrac{1}{ x+1}\hspace{1.5em}\\~\\ \implies x = 1 + \cfrac{1}{ x+1}\hspace{1.5em}\\~\\ \end{align}}\)

OpenStudy (anonymous):

So the value of x =2?

OpenStudy (mathmath333):

no u will get the quadratic equation , which will give two values

OpenStudy (anonymous):

Yes, sorry.. x=1 is what I get

OpenStudy (anonymous):

@ganishe8 this is fast, real fast would like to know who taught you maths

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} x +1= 2 + \dfrac{1}{ x+1}\hspace{1.5em}\\~\\ ( x +1)^2= 2(x+1) + 1\hspace{1.5em}\\~\\ ( x +1)^2-2(x+1)+1= 2\hspace{1.5em}\\~\\ \left(( x +1)-1\right)^2= 2\hspace{1.5em}\\~\\ x=\pm\sqrt 2 \end{align}}\)

OpenStudy (anonymous):

hmmm, yes thanks mathmath333

OpenStudy (mathmath333):

note that here only \(x=+\sqrt2\) is possible and not \(x\neq -\sqrt2\)

OpenStudy (anonymous):

so the ans is \[\sqrt{2}\] ?

OpenStudy (mathmath333):

yes